# Hyperparameter Optimization in Machine Learning

### v2

@@ 8.1 Optimizing for multiple objectives
In many real world settings there is not just one, but multiple, potentially conflicting, objectives of interest. For example, when working with hardware-constrained systems one desires predictive models that are accurate and have low resource consumption. Assume one needs to decide between two solutions $\lambda_{1},\lambda_{2}$. Each individual objective induces a distinct ordinal relationship between the candidates, but while $\lambda_{1}$ might dominate $\lambda_{2}$ in one objective, it might be worse in another. Because there is rarely a single best solution, the multi-objective hyperparameter optimization (MO) problem is to identify the Pareto front—the set of all non-dominated solutions—to help users make an informed trade-off decision. More formally, consider a function $f:\Lambda\to\mathcal{Y}$ mapping points in domain $\Lambda$ to a space $\mathcal{Y}$, which in the following we consider to be an $n$-dimensional objective space that we aim to minimize. For two points $\lambda_{1},\lambda_{2}\in\Lambda$, we say $\lambda_{1}$ weakly dominates $\lambda_{2}$ if $f(\lambda_{1})_{i}\leq f(\lambda_{2})_{i}$ for all objectives $i\in[n]$. We write $\lambda_{1}\succeq\lambda_{2}$ to denote weak domination. Furthermore, $\lambda_{1}$ is said to dominate $\lambda_{2}$ if $\lambda_{1}\succeq\lambda_{2}$ and $f(\lambda_{1})_{i}<f(\lambda_{2})_{i}$ for some objective $i$. We write $\lambda_{1}\succ\lambda_{2}$ for domination. The Pareto front $\mathcal{P}_{f}$ contains all nondominated points in $\Lambda$, formally:
$\mathcal{P}_{f}=\{\lambda\in\Lambda\,|\,\nexists\lambda^{\prime}\in\Lambda:\lambda^{\prime}\succ\lambda\}.$
Since the Pareto front is often infinite, multi-objective optimizers seek a finite approximation set $B\subset\mathcal{Y}$ of objective vectors near the Pareto front. The quality of $B$ can be measured by the dominated hypervolume indicator $\mathcal{H}$ [Zitzler1998:Multiobjective]. Given a reference point $\mathbf{r}$, $\mathcal{H}(B)$ equals the hypervolume dominated by $B$:
$\mathcal{H}(B)=\text{Vol}\left(\{\mathbf{y}\in\mathcal{Y}=\mathbb{R}^{n}\,|\,\exists\mathbf{y}^{\prime}\in B:\mathbf{y}^{\prime}\succeq\mathbf{y}\wedge\mathbf{y}\succeq\mathbf{r}\}\right).$
@@
Examples
A popular example of an MO scenario is the case of algorithmic fairness which aims to train accurate models subject to fairness criteria. Prior work in the field has proposed several fairness criteria, whose suitability depends on the application domain, and it has been shown that criteria, commonly referred to as acceptance-based and error-based, can conflict with each other [friedler2016possibility, fairness2018verma, kleinberg2018inherent]. However, obtaining solutions that satisfy various definitions to some extent and expose empirical trade-offs between various definitions can still be important from societal perspectives. Another way to tackle MO scenarios is the case of gradient-free optimization with stochastic constraints. For example, when enhancing the accuracy of a machine learning model, considerations may include constraints on model memory or prediction latency. This occurs in the practical application of deploying machine learning models on mobile devices, where memory cannot exceed a certain pre-specified threshold, or when models power devices with tight requirements on inference latency. The objective and constraint functions in such cases are real-valued and observed together. Alternatively, in tuning a deep neural network (DNN), one may aim to prevent training failures due to out-of-memory errors. Here, the constraint feedback is binary, indicating whether the constraint is violated, and the objective is not observed under constraint violation. It is common to treat constraints with equal importance to objectives, employing a surrogate model as discussed in Chapter 4. Unfeasible evaluations incur a cost, such as wasted resources or compute node failure, necessitating minimization of their occurrence. However, avoiding the unfeasible region should not compromise convergence to optimal hyperparameters. Often, optimal configurations lie at the boundary of the feasible region, presenting a challenge in balancing the costs of unfeasible evaluations and the need for exploration.
@@ 8.1.1 Multi-objective optimization
Multi-objective Bayesian optimization (MBO) techniques can be categorized into three classes. First, scalarization-based MBO methods map the vector of all objectives to a scalar [Knowles2006:Parego, Zhang2009:Expensive, Nakayama2009:Sequential, Paria2019:Flexible, Golovin2020:Random] and then use conventional single-objective BO techniques. The scalarization is defined by using a weight vector $\mathbf{w}\in\mathbb{R}^{n}$.
Some common choices for $V$ are:
  • Inner product between objective and weight vectors (also known as random weights): $V_{\text{RW}}(f(\lambda),\mathbf{w})=f(\lambda)^{\top}\mathbf{w}$;
@@
    $V_{\text{ParEGO}}(f(\lambda),\mathbf{w})=\max_{j\in\{1,\dots,n\}}(w_{j}f(\lambda)_{j})+\rho(f(\lambda)^{\top}\mathbf{w});$
@@
    $V_{\text{Golovin}}(f(\lambda),\mathbf{w})=\min_{j\in\{1,\dots,n\}}(\max(0,f(\lambda)_{j}/w_{j}))^{n}.$
Note that by optimizing different scalarizations, solutions along the Pareto front can be obtained. The choice of $V$ influences the subset of the Pareto front that is found, and the right selection might change depending on the task at hand.
@@ 8.1.2 Constrained hyperparameter optimization
Formally, the goal in a constrained optimization scenario is to minimize the target response function $f(\lambda)$, subject to a constraint $c(\lambda)\leq\delta$. Here, we limit our attention to modeling the feasible region by a single function $c(x)$. As the latent constraints are pairwise conditionally independent, an extension to multiple constraints is straightforward, yet notationally cumbersome. Both $f(\lambda):\Lambda\rightarrow\mathbb{R}$ and $c(x):\Lambda\rightarrow\mathbb{R}$ are unknown and need to be queried sequentially. The constrained optimization problem is defined as follows:
$\min_{\{\lambda\in\Lambda\,:\,c(\lambda)\leq\delta\}}f(\lambda),$ (8.1)
@@
where $\delta\in\mathbb{R}$ is a threshold on the constraint. For instance, in the case of BO, the latent functions $f(\lambda)$ and $c(\lambda)$ can be conditionally independent in our surrogate model, with different GP priors placed on them.
@@
We consider two different setups, depending on what information is observed about the constraint. Many existing works [Gardner14, Gelbart14, Lobato15a] assumes that real-valued feedback is obtained on $c(\lambda)$, just as for $f(\lambda)$. In this case, both latent function can be represented as GPs with Gaussian noise. Unfortunately, this setup does not cover practically important use cases of constrained hyperparameter optimization. For example, if training a DNN fails with an out-of-memory error, we cannot observe the amount of memory requested just before the crash. Rather, we only know that an out-of-memory error occurred, which is a binary outcome. Covering such use cases requires handling binary feedback on $c(x)$, even though this is technically more difficult. We can assume an evaluation returns $z_{f}\sim\mathcal{N}(z_{f}|f(\lambda),\alpha_{f}^{-1})$ and $z_{c}\in\{-1,+1\}$, where $z_{c}=-1$ for a feasible, $z_{c}=+1$ for an unfeasible point. Here, $z_{f}$ is the observed value of the objective function, while $z_{c}$ is the observed value of the constraint. We never observe the latent constraint function $c(\lambda)$ directly. We assume $z_{c}\sim\sigma(z_{c}c(\lambda))$, where $\sigma(t)=\frac{1}{1+e^{-t}}$ is the logistic sigmoid, but other choices are possible. We can then rewrite the constrained optimization problem (8.1) as follows:
$y_{\star}=\min_{\lambda\in\Lambda}\left\{y(\lambda)\;\|\;P(z_{c}=+1|\lambda)=\sigma(c(\lambda))\leq\sigma(\delta)\right\}.$
@@
This formulation is similar to the one proposed in [Gelbart14]. The parameter $\sigma(\delta)\in(0,1)$ controls the size of the (random) feasible region for defining $y_{\star}$. Finally, note that in the example of out of memory errors training failures, the criterion observation $z_{y}$ is obtained only if $z_{c}=-1$: if a training run crashes, a validation error is not obtained for the queried configuration. Apart from an experiment in [Gelbart14], the case of contrained BO work covering the binary feedback case is covered in [perrone_cmes].
@@
The most established technique to tackle constrained BO is constrained EI (cEI) [Gardner14, Gelbart14, snoek_scalable_2015, letham2019]. A separate regression model is used to learn the constraint function $c(\lambda)$ (typically a GP), and EI is modified in two ways. First, the expected amount of improvement of an evaluation is computed only with respect to the current feasible minimum. Second, hyperparameters with a large probability of satisfying the constraint are encouraged by optimizing $cEI(\lambda)=\mathbb{P}\{c(\lambda)\leq\delta\}EI(\lambda)$, where $\mathbb{P}\{c(\lambda)\leq\delta\}$ is the posterior probability of $\lambda$ being feasible under the constraint model, and $EI(\lambda)$ is the standard EI acquisition function, see Section 4.2.3.
@@
Several issues with cEI are detailed in [Lobato15a]. First, the current feasible minimum has to be known, which is problematic if all initial evaluations are unfeasible. A workaround is to use a different acquisition function, initially focused on finding a feasible point [Gelbart14]. In addition, the probability of constraint violation is not explicitly accounted for in cEI. A confidence parameter features in [Gelbart14], but is only used to recommend the final hyperparameter configuration. Another approach was proposed in [Lobato15a], where PES is extended to the constrained case. Constrained PES (cPES) can outperform cEI and does not require the workarounds mentioned above. However, it is complex to implement, expensive to evaluate, and unsuitable for binary constraint feedback. Drawing from numerical optimization, different generalizations of EI to the constrained case are developed in [Picheny2016] and [Ariafar2019]. The authors represent the constrained minimum by way of Lagrange multipliers, and the resulting query selection problem is solved as a sequence of unconstrained problems. An alternative option that does not require expensive computations is the extension of another acquisition function to the constrained case, namely Max-value Entropy Search [perrone_cmes]. This acquisition function handles binary constraint feedback, does not require numerical quadrature [Picheny2016], and does not come with a large set of extra hyperparameters [Ariafar2019].
@@ 8.2 Hyperparameter schedules and online HPO
The vast majority of learning algorithms do not produce an hypothesis in one shot, but rather refine it incrementally over the course of many iterations. In the previous chapters we already discussed algorithms designed around this observation, such as those leveraging multi-fidelity approaches (Chapter 5), and hypergradients (Section 7.3), as well as some evolutionary algorithms like population-based training (Section 6.3). When this is the case, it may be sensible to consider hyperparameter schedules rather than fixed values. One of the most prominent example in the literature are the ubiquitous learning rate schedules when training neural nets [bengio2012practical, loshchilov2017sgdr, goyal2017accurate]. In this section we overview such setup and briefly discuss how one may tackle the HPO problem as an instance of online optimization.
Mimicking the notation of Section 7.3, let us consider an algorithm $A:\Lambda\to\mathcal{H}$ as a composition of iterates $\Phi_{t}:\mathcal{H}\times\Lambda\times\rightarrow\mathcal{H}$. 111 We disregard the dependence of the maps on data, for notational simplicity. Such transitions maps need not be necessarily differentiable, although several methods that exploit the iterative view rely on this assumption. This time, we let the hyperparameter configuration also vary as a function of the step. That is, given an initial model $h_{0}\in\mathcal{H}$:
$A(\lambda)=\Phi_{T}(\cdot,\lambda_{T})\circ\Phi_{T-1}(\cdot,\lambda_{T-1})\circ\dots\circ\Phi_{1}(h_{0},\lambda_{1}).$ (8.2)
for some $T>0$, where we generally view $\lambda\in\Lambda^{T}$ as an hyperparameter sequence. The iterations represented by the iterates $\Phi_{t}$ can correspond to different granularities or spans of operations within the training process, such as single stochastic gradient descent steps, entire epochs, or any other sequence of operations.
Hyperparameter iterates may be left “free to float”, in which case one can formulate a related HPO problem following the general setup of Section 2.2, but this time searching for configurations in $\Lambda^{T}$ rather than in $\Lambda$. This is seldom done in practice due to the complications arising from an expanded search space, few exceptions being [maclaurin2015gradient, wu2018learning]. More commonly, the iterates are expressed as a function of the step $t$ and a set of other fixed hyperparameters $\lambda^{\prime}\in\Lambda^{\prime}$ Learning rate schedules are a primary example. For instance, a triangular schedule (i.e. warmup followed by linear decay) is given by:
@@
However, perhaps one main advantage of explicitly taking an iterative view is that it enable recasting the HPO problem as an instance of online optimization [shalev2012online], albeit a non-standard one. Online HPO techniques attempt to speed up the HPO procedure by finding a performing hyperparamter schedule in one go, during a single execution of the learning algorithm. A key difficulty is that one does not have access to the true objective $\mathscr{M}$. Indeed, being able to compute (an unbiased estimate of) $\mathscr{M}$ requires terminating the learning algorithm, which would defeat the purpose of online HPO. In this sense, these techniques do not – or rather, cannot – strictly solve Problem (2.4)-(2.5), but rather rely on approximations, heuristics or surrogate objectives. This is akin to what we discussed in Chapter 5, however multi-fidelity techniques only use early observations to take decisions or fit models of the response surface, and do not change hyperparameters during execution. Examples of online HPO include [orabona2016coin, baydin_2017_online, luketina2016scalable, jaderberg2017population, lorenzo2017particle, mackay2019self, donini2019opt, vicol2021unbiased, micaelli2021gradient, chandra2022gradient], with gradient-based techniques, and especially forward iterative differentiation, being proposed by various authors.
Another stream of recent research focuses on tuning hyperparameters of online learning algorithms, such as contextual bandit [li2010contextual] methods like UCB [auer2002finite] or Thomson Sampling [agrawal2013thompson]. In these algorithms, hyperparameters affect the agent’s exploration-exploitation trade-off and must adapt in real time. bouneffouf2020hyper treat this as a nested bandit problem, tuning exploration parameters online via meta-bandits. ding2022syndicated generalize this idea with Syndicated Bandits, a multi-layer optimization scheme that tunes several hyperparameters simultaneously, with favorable regret guarantees. kang2024online further extend this line by framing tuning as a continuum-armed bandit problem, enabling efficient search in continuous spaces. These methods depart from gradient-based HPO discussed above by tightly coupling hyperparameter adaptation with decision-time feedback, offering principled and scalable alternatives tailored to the online nature of bandit learning.
Online HPO remains an area of active research.

### v3

@@ 8.1 Optimizing for Multiple Objectives
In many real world settings there is not just one, but multiple, potentially conflicting, objectives of interest. For example, when working with hardware-constrained systems one desires predictive models that are accurate and have low resource consumption. Assume one needs to decide between two solutions $\lambda_{1},\lambda_{2}$. Each individual objective induces a distinct ordinal relationship between the candidates, but while $\lambda_{1}$ might dominate $\lambda_{2}$ in one objective, it might be worse in another. Because there is rarely a single best solution, the multi-objective hyperparameter optimization (MO) problem is to identify the Pareto front — the set of all non-dominated solutions — to help users make an informed trade-off decision. More formally, consider a function $f\mathrel{\mathop{\ordinarycolon}}\Lambda\to\mathcal{Y}$ mapping points in domain $\Lambda$ to a space $\mathcal{Y}$, which in the following we consider to be an $n$-dimensional objective space that we aim to minimize. For two points $\lambda_{1},\lambda_{2}\in\Lambda$, we say $\lambda_{1}$ weakly dominates $\lambda_{2}$ if $f(\lambda_{1})_{i}\leq f(\lambda_{2})_{i}$ for all objectives $i\in[n]$. We write $\lambda_{1}\succeq\lambda_{2}$ to denote weak domination. Furthermore, $\lambda_{1}$ is said to dominate $\lambda_{2}$ if $\lambda_{1}\succeq\lambda_{2}$ and $f(\lambda_{1})_{i}<f(\lambda_{2})_{i}$ for some objective $i$. We write $\lambda_{1}\succ\lambda_{2}$ for domination. The Pareto front $\mathcal{P}_{f}$ contains all nondominated points in $\Lambda$, formally:
$\mathcal{P}_{f}=\{\lambda\in\Lambda\mid\nexists\lambda^{\prime}\in\Lambda\mathrel{\mathop{\ordinarycolon}}\lambda^{\prime}\succ\lambda\}.$
Since the Pareto front is often infinite, multi-objective optimizers seek a finite approximation set $B\subset\mathcal{Y}$ of objective vectors near the Pareto front. The quality of $B$ can be measured by the dominated hypervolume indicator $\mathcal{H}$ [Zitzler1998:Multiobjective]. Given a reference point $r$, $\mathcal{H}(B)$ equals the hypervolume dominated by $B$:
$\mathcal{H}(B)=\text{Vol}\left(\{y\in\mathcal{Y}\subseteq\mathds{R}^{n}\mid\exists y^{\prime}\in B\mathrel{\mathop{\ordinarycolon}}y^{\prime}\succeq y\wedge y\succeq r\}\right).$
@@
Examples.
A popular example of an MO scenario is the case of algorithmic fairness which aims to train accurate models subject to fairness criteria. Prior work in the field has proposed several fairness criteria, whose suitability depends on the application domain, and it has been shown that criteria, commonly referred to as acceptance-based and error-based, can conflict with each other [friedler2016possibility, fairness2018verma, kleinberg2018inherent]. However, obtaining solutions that satisfy various definitions to some extent and expose empirical trade-offs between various definitions can still be important from a societal perspective. Another way to tackle MO scenarios is the case of gradient-free optimization with stochastic constraints. For example, when enhancing the accuracy of a machine learning model, considerations may include constraints on model memory or prediction latency. This occurs in the practical application of deploying machine learning models on mobile devices, where memory cannot exceed a certain pre-specified threshold, or when models power devices with tight requirements on inference latency. The objective and constraint functions in such cases are real-valued and observed together. Alternatively, in tuning a deep neural network (DNN), one may aim to prevent training failures due to out-of-memory errors. Here, the constraint feedback is binary, indicating whether the constraint is violated, and the objective is not observed under constraint violation. It is common to treat constraints with equal importance to objectives, employing a surrogate model as discussed in Chapter 4. Unfeasible evaluations incur a cost, such as wasted resources or compute node failure, necessitating minimization of their occurrence. However, avoiding the unfeasible region should not compromise convergence to optimal hyperparameters. Often, optimal configurations lie at the boundary of the feasible region, presenting a challenge in balancing the costs of unfeasible evaluations and the need for exploration.
@@ 8.1.1 Multi-Objective Optimization
Multi-objective Bayesian optimization (MBO) techniques can be categorized into three classes. First, scalarization-based MBO methods map the vector of all objectives to a scalar $V$ [Knowles2006:Parego, Zhang2009:Expensive, Nakayama2009:Sequential, Paria2019:Flexible, Golovin2020:Random] and then use conventional single-objective BO techniques. The scalarization is defined by using a weight vector $w\in\mathds{R}^{n}$. Some common choices for $V$ are:
  • Inner product between objective and weight vectors (also known as random weights): $V_{\text{RW}}(f(\lambda),w)=f(\lambda)^{\top}w$;
@@
    $V_{\text{ParEGO}}(f(\lambda),w)=\max_{j\in\{1,\dots,n\}}(w_{j}f(\lambda)_{j})+\rho(f(\lambda)^{\top}w);$
@@
    $V_{\text{Golovin}}(f(\lambda),w)=\min_{j\in\{1,\dots,n\}}(\max(0,f(\lambda)_{j}/w_{j}))^{n}.$
Note that by optimizing different scalarizations, different solutions along the Pareto front can be obtained. The choice of $V$ influences the subset of the Pareto front that is found, and the right selection might change depending on the task at hand.
@@ 8.1.2 Constrained Hyperparameter Optimization
Formally, the goal in a constrained optimization scenario is to minimize the target response function $f(\lambda)$, subject to a constraint $c(\lambda)\leq\delta$. Here, we limit our attention to modeling the feasible region by a single function $c(x)$. As the latent constraints are pairwise conditionally independent, an extension to multiple constraints is straightforward, yet notationally cumbersome. Both $f(\lambda)\mathrel{\mathop{\ordinarycolon}}\Lambda\rightarrow\mathds{R}$ and $c(x)\mathrel{\mathop{\ordinarycolon}}\Lambda\rightarrow\mathds{R}$ are unknown and need to be queried sequentially. The constrained optimization problem is defined as follows:
$\min_{\lambda\in\Lambda}f(\lambda),\quad\text{s.t. }\;c(\lambda)\leq\delta$ (8.1)
@@
where $\delta\in\mathds{R}$ is a threshold on the constraint. For instance, in the case of BO, the latent functions $f(\lambda)$ and $c(\lambda)$ can be conditionally independent in our surrogate model, with different Gaussian process priors placed on them.
@@
We consider two different setups, depending on what information is observed about the constraint. Many existing works [Gardner14, Gelbart14, Lobato15a] assume that real-valued feedback is obtained on $c(\lambda)$, just as for $f(\lambda)$. In this case, both latent functions can be represented as Gaussian processes with Gaussian noise. Unfortunately, this setup does not cover practically important use cases of constrained hyperparameter optimization. For example, if training a deep network fails with an out-of-memory error, we cannot observe the amount of memory requested just before the crash. Rather, we only know that an out-of-memory error occurred, which is a binary outcome. Covering such use cases requires handling binary feedback on $c(x)$, even though this is technically more difficult. We can assume an evaluation returns $z_{f}\sim\mathcal{N}(z_{f}|f(\lambda),\alpha_{f}^{-1})$ and $z_{c}\in\{-1,+1\}$, where $z_{c}=-1$ for a feasible, $z_{c}=+1$ for an unfeasible point. Here, $z_{f}$ is the observed value of the objective function, while $z_{c}$ is the observed value of the constraint. We never observe the latent constraint function $c(\lambda)$ directly. We assume $z_{c}\sim\sigma(z_{c}c(\lambda))$, where $\sigma(t)=\frac{1}{1+e^{-t}}$ is the logistic sigmoid, but other choices are possible. We can then rewrite the constrained optimization problem (8.1) as follows:
$y_{\star}=\min_{\lambda\in\Lambda}\left\{f(\lambda)\mid\mathbb{P}[z_{c}=1|\lambda]=\sigma(c(\lambda))\leq\sigma(\delta)\right\}.$
@@
This formulation is similar to the one proposed by Gelbart14. The parameter $\sigma(\delta)\in(0,1)$ controls the size of the (random) feasible region for defining $y_{\star}$. Finally, note that in the example of out of memory errors training failures, the criterion observation $z_{y}$ is obtained only if $z_{c}=-1$: if a training run crashes, a validation error is not obtained for the queried configuration. Apart from an experiment by Gelbart14, the case of constrained BO work covering the binary feedback case is covered by perrone_cmes.
@@
The most established technique to tackle constrained BO is constrained EI (cEI) [Gardner14, Gelbart14, snoek_scalable_2015, letham2019]. A separate regression model is used to learn the constraint function $c(\lambda)$, typically a Gaussian process, and EI is modified in two ways. First, the expected amount of improvement of an evaluation is computed only with respect to the current feasible minimum. Second, hyperparameters with a large probability of satisfying the constraint are encouraged by optimizing $cEI(\lambda)=\mathbb{P}[c(\lambda)\leq\delta]\cdot EI(\lambda)$, where $\mathbb{P}[c(\lambda)\leq\delta]$ is the posterior probability of $\lambda$ being feasible under the constraint model, and $EI(\lambda)$ is the standard EI acquisition function, see Section 4.2.3.
@@
Several issues with cEI have been noted by Lobato15a. First, the current feasible minimum has to be known, which is problematic if all initial evaluations are unfeasible. A workaround is to use a different acquisition function, initially focused on finding a feasible point [Gelbart14]. In addition, the probability of constraint violation is not explicitly accounted for in cEI. A confidence parameter was introduced by Gelbart14, but it is only used to recommend the final hyperparameter configuration. Another approach was proposed by Lobato15a, where PES is extended to the constrained case. Constrained PES (cPES) can outperform cEI and does not require the workarounds mentioned above. However, it is complex to implement, expensive to evaluate, and unsuitable for binary constraint feedback. Drawing from numerical optimization, different generalizations of EI to the constrained case were developed by Picheny2016 and by Ariafar2019. The authors represent the constrained minimum by way of Lagrange multipliers, and the resulting query selection problem is solved as a sequence of unconstrained problems. An alternative option that does not require expensive computations is the extension of another acquisition function to the constrained case, namely Max-value Entropy Search [perrone_cmes]. This acquisition function handles binary constraint feedback, does not require numerical quadrature [Picheny2016], and does not come with a large set of extra hyperparameters [Ariafar2019].
@@ 8.2 Hyperparameter Schedules and Online HPO
The vast majority of learning algorithms do not produce a hypothesis in one shot, but rather refine it incrementally over the course of many iterations. In the previous chapters we already discussed algorithms designed around this observation, such as those leveraging multi-fidelity approaches (Chapter 5), and hypergradients (Section 7.3), as well as some evolutionary algorithms like population-based training (Section 6.3). When this is the case, it may be sensible to consider hyperparameter schedules rather than fixed values. One of the most prominent examples in the literature are the ubiquitous learning rate schedules when training neural nets [bengio2012practical, loshchilov2017sgdr, goyal2017accurate]. In this section we overview such setup and briefly discuss how one may tackle the HPO problem as an instance of online optimization.
Mimicking the notation of Section 7.3, let us consider an algorithm $A\mathrel{\mathop{\ordinarycolon}}\Lambda\to\mathcal{H}$ as a composition of iterates $\Phi_{t}\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\times\Lambda\rightarrow\mathcal{H}$.  [^1] We disregard the dependence of the maps on data, for notational simplicity. Such transitions maps need not be necessarily differentiable, although several methods that exploit the iterative view rely on this assumption. This time, we let the hyperparameter configuration also vary as a function of the step. That is, given an initial model $h_{0}\in\mathcal{H}$:
$A(\lambda)=\Phi_{T}(\cdot,\lambda_{T})\circ\Phi_{T-1}(\cdot,\lambda_{T-1})\circ\dots\circ\Phi_{1}(h_{0},\lambda_{1}),$ (8.2)
for some $T>0$, where we view $\lambda\in\Lambda^{T}$ as a hyperparameter sequence. The iterations represented by the iterates $\Phi_{t}$ can correspond to different granularities or spans of operations within the training process, such as single stochastic gradient descent steps, entire epochs, or any other sequence of operations.
Hyperparameter iterates may be left “free to float”, in which case one can formulate a related HPO problem following the general setup of Section 2.2, but this time searching for configurations in $\Lambda^{T}$ rather than in $\Lambda$. This is seldom done in practice due to the complications arising from an expanded search space, few exceptions being [maclaurin2015gradient, wu2018learning]. More commonly, the iterates are expressed as a function of the step $t$ and a set of other fixed hyperparameters $\lambda^{\prime}\in\Lambda^{\prime}$. Learning rate schedules are a primary example. For instance, a triangular schedule (i.e. warmup followed by linear decay) is given by:
@@
One main advantage of explicitly taking an iterative view is that it enable recasting the HPO problem as an instance of online optimization [shalev2012online], albeit a non-standard one. Online HPO techniques attempt to speed up the HPO procedure by finding a performing hyperparameter schedule in one go, during a single execution of the learning algorithm. A key difficulty is that one does not have access to the true objective $\mathscr{M}$. Indeed, being able to compute (an unbiased estimate of) $\mathscr{M}$ requires terminating the learning algorithm, which would defeat the purpose of online HPO. In this sense, these techniques do not – or rather, cannot – strictly solve Problem (2.4)-(2.5), but rather rely on approximations, heuristics or surrogate objectives. This is akin to what we discussed in Chapter 5, however multi-fidelity techniques only use early observations to take decisions or fit models of the response surface, and do not change hyperparameters during execution. Examples of online HPO include [orabona2016coin, baydin_2017_online, luketina2016scalable, jaderberg2017population, lorenzo2017particle, mackay2019self, donini2019opt, vicol2021unbiased, micaelli2021gradient, chandra2022gradient], with gradient-based techniques, and especially forward iterative differentiation, being proposed by various authors.
Another stream of recent research focuses on tuning hyperparameters of online learning algorithms, such as contextual bandit [li2010contextual] methods like UCB [auer2002finite] or Thomson Sampling [agrawal2013thompson]. In these algorithms, hyperparameters affect the agent’s exploration-exploitation trade-off and must adapt in real time. bouneffouf2020hyper treat this as a nested bandit problem, tuning exploration parameters online via meta-bandits. ding2022syndicated generalize this idea with Syndicated Bandits, a multi-layer optimization scheme that tunes several hyperparameters simultaneously, with favorable regret guarantees. kang2024online further extend this line by framing tuning as a continuum-armed bandit problem, enabling efficient search in continuous spaces. These methods depart from gradient-based HPO discussed above by tightly coupling hyperparameter adaptation with decision-time feedback, offering principled and scalable alternatives tailored to the online nature of bandit learning. Online HPO remains an area of active research.